If the coefficient of static friction is 0.35 and the normal force is 80 newtons, what is the maximum frictional force of the surface acting on the object?
1
Expert's answer
2012-05-22T10:48:52-0400
Fmax = N * k = 80 * 0.35 = 28 newtons.